OpenStudy (vshiroky):

Find the slope intercept for 3x-8y=1 I think the slope is -3/8 and y intercept is (0,-1/8)

jimthompson5910 (jim_thompson5910):

close, but the slope is positive

OpenStudy (vshiroky):

Is that because 3 is - and 8 is -

jimthompson5910 (jim_thompson5910):

3x-8y=1 -8y=1-3x -8y = -3x + 1 y = (-3x + 1)/(-8) y = (-3x)/(-8) + 1/(-8) y = (3/8)x - 1/8 ... the two negatives cancel to make a positive

OpenStudy (vshiroky):

Ok then for something like (-3,0) and (0,-8) The slope intercept would be -8/3 because only one is negative?

OpenStudy (johnweldon1993):

The SLOPE would be that yes! Not the intercept

OpenStudy (vshiroky):

Lol that's what I meant

OpenStudy (johnweldon1993):

Okay good :) lol